OpenStudy (anonymous):

HELP FOR A MEDAL AND TO HAVE A FAN ! What is the distance from (-2, -4) to (1, 0)? 3 units 4 units 5 units 6 units

OpenStudy (sw050399):

You have to use the distance formula. \[\sqrt{(x _{1}-x _{2})^2-(y _{2}-y _{1})^2}\]

OpenStudy (sw050399):

Just plug in your points. It doesn't matter which one is (x1, y1) or (x2, y2), just make sure they correspond.

OpenStudy (anonymous):

You could use the distance formula like sw said, or you can understand where it came from first. It is derived from the length of the hypotenuse formula in triangles, since you can actually just draw a triangle with those points and link them directly with the hypotenuse.

OpenStudy (anonymous):

Thanks guys I think I got it (:

OpenStudy (nurali):

distance betweeen X values = -3 distance betweeen Y values = -4 using Pythagoras theorem, X^2 + Y^2 = distance between points^2 in other words Y^2 (-3)^2 +(-4)^2 =25 Sq root of 25 = 5 distance between points =5 units

OpenStudy (sw050399):

@Nurali is correct. The formula is adding, not subtracting as I accidentally put.